OpenStudy (anonymous):

hi, how do you isolate x in this equation: a/b = (.5) exponent x?

OpenStudy (anonymous):

by exponent do you mean \[0.5^x\]or \[exp(x)\]where exp(x) is equivalent to the exponential function \[e^x\]

OpenStudy (anonymous):

either way, use logarithms

OpenStudy (anonymous):

\[x = -\frac{\ln(\frac{a}{b})}{\ln(2)}\]

OpenStudy (anonymous):

the first one...

OpenStudy (anonymous):

If it's the first one then my above answer stands. You can also write it as \[x = \frac{\ln(b) - \ln(a)}{\ln(2)}\]

OpenStudy (anonymous):

ok thanks!